OpenStudy (anonymous):

How do you find all the solutions for x in the equation (x^2+6)^2/3=16

myininaya (myininaya):

raise both sides to the 3/2 power

myininaya (myininaya):

you will get \[x^2+6=16^\frac{3}{2}\]

myininaya (myininaya):

square root of 16 is 4 4^3=64 so we have \[x^2+6=64\]

myininaya (myininaya):

i will leave the rest to you

myininaya (myininaya):

let me know if you run into trouble

myininaya (myininaya):

actually i think we can say plus or minus

myininaya (myininaya):

\[x^2+6=\pm 64\]

myininaya (myininaya):

but you should check both solutions to make sure they are actually solutions

OpenStudy (anonymous):

\[(x^{2}+6)^{2/3}=16\] this is the actual equation, sorry

myininaya (myininaya):

ok that is what i thought it was lol